OpenStudy (anonymous):

Help please!! Find the limit of the function by using direct substitution. limit as x approaches quantity pi divided by two of quantity three times e to the x times cosine of x.

OpenStudy (loser66):

Translate the words into math notation, please

OpenStudy (anonymous):

\[\lim(3e^x \cos x)_{x \rightarrow \pi/2}\]

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

\[\lim_{x\rightarrow \pi/2} (3e^x cos x)\]

OpenStudy (loser66):

so, where you see x, put pi/2. Actually cos (pi/2) =0, hence the whole thing =0

OpenStudy (anonymous):

are you kidding me the whole thing is 0??? -.-

OpenStudy (loser66):

yup

OpenStudy (anonymous):

insert crying laughing emoji here. thanks bruh.